Uan Newtons method to calculator x for f(x) = 2x + 8 pven that x = 2 x2=1.00 *2=125 X: 050
Answer #1

Solution - Here fox) = 2x² +8, 20=2 fa) = 6x² .. By Newtons Method we have x = xo-fa) - 2- (2x2²+8) 6x2² - 2- 24 24 =2-1=1.